what is the maxium y value for f(x)=x^2+6x+8
Answer by aquyu214(10)

f(x)=x^2+2*3*x+3^2-3^2+8
=x^2+2*3*x+3^2-1
=(X+3)^2-1
when X=-3 get the minimum value -1
X changed then Y can be infinite, there is no maximum y value